Reports 1-1 of 1 Clear search Modify search
Detector Operation (Alignment Values)
berni - 15:35 Monday 24 November 2025 (68228) Print this report
Operator Report - Morning shift

ITF found in Troubleshooting mode with the following issues:

  • BS loops opened, problem know: 68224.
  • PR vertical correction saturating; to be recovered by the experts.
  • SIB1 IP open; likely because of the previous problem.
  • SPRB and SWEB vertical correction to be recovered.
  • temperature in NE and WE outside the thresholds; to be investigated.

At around 8:00 UTC I recovered SIB1, SPRB and SWEB: in parallel Davide went to the END buildings to adjust the set points of the boilers.

At around 10:00 UTC the SAT team went in CB to work on the BS.

At around 10:30 UTC Paolo recovered the PR suspsension: after that I recovered the aligment of the north cavity and I was able to lock it on the IR.

At around 10:30 UTC Romain started to perform some noise injection on the DET bench.

At around 10:45 UTC the SAT team recovered the BS suspension; after that I manually realigned the WEST cavity and I was able to relock the ITF on the IR restoring the automation.

The BS suspension failed again at around 11:40 UTC;  situation in progress.

 

Software

BacNet restarted at 10:02 UTC because it was providing flat data.

 

SBE

SNEB loop closed and vertical position with step motors at 10:15 UTC

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×